A 50ve change flow through a conductor if the p.d between it end 20mv find the work done by the change 


1
Expert's answer
2021-02-19T18:58:23-0500

By definition, the work done by charge is given as:


"W = -Uq"

where "U = 20mV =0.02V" is the potential difference, and


"q = 50ve = 50\cdot 1.6\times 10^{-19}C = 8\times 10^{-18}C"

is the charge ("1.6\times 10^{-19}C" is the elementary charge). Thus, obtain:

"W = -0.02V\cdot 8\times 10^{-18}C = -1.6\times 10^{-19}J"

Answer. "-1.6\times 10^{-19}J".